# TimeLDM: Latent Diffusion Model for Unconditional Time Series Generation

### v1

@@ 3.4 Training and Sampling
4 Experiments
In this section, we empirically examine the capability of our model for time series generation tasks. We first introduce our experimental setup in Section 4.1. Then, we report and analyze the evaluation main results in Section 4.2. Moreover, we present a further study of TimeLDM on different lengths of time series. Finally, we perform the ablation study of TimeLDM in Section 4.3.
4.1 Experimental Setups
Datasets. As Table 1 shows, we utilize four different datasets to evaluate our model. The datasets include: Sine is a simulated dataset, we generate 10000 samples with 5 features in sinusoidal sequence, and each feature has independent frequencies and phases; Stocks is the Google stock price information from 2004 to 2019, this stock data is presented daily information and includes 6 features. ETTh is built from electricity transformers on 15 minutes basis, including load and oil temperature from July 2016 to July 2018; MuJoCo is the multivariate physics simulation time series data with 14 features.
Table 1: The details of experimental datasets, where the Sines and Mujoco are simulated datasets, the Stocks and ETTh are realistic datasets.
Baseline. We compare our TimeLDM against six unconditional time series generation methods, including: Diffusion-TS [4] (ICLR2024), DiffTime [3] (NeurIPS2023), and Diffwave [11] (ICLR2021), which are based on diffusion architectures. Furthermore, we select TimeGAN [13] (NeurIPS2019), which relies on GAN-based models as same as Cot-GAN [17] (NeurIPS2020). Lastly, we consider TimeVAE [19] (ICLR2022), which is built on a VAE-based approach.
Metric Methods Sines Stocks ETTh MuJoCo
TimeLDM 0.004$\pm$.001 0.032$\pm$.007 0.034$\pm$.003 0.012$\pm$.001
Diffusion-TS 0.006$\pm$.000 0.147$\pm$.025 0.116$\pm$.010 0.013$\pm$.001
Context-FID TimeGAN 0.101$\pm$.014 0.103$\pm$.013 0.300$\pm$.013 0.563$\pm$.052
Score TimeVAE 0.307$\pm$.060 0.215$\pm$.035 0.805$\pm$.186 0.251$\pm$.015
Diffwave 0.014$\pm$.002 0.232$\pm$.032 0.873$\pm$.061 0.393$\pm$.041
DiffTime 0.006$\pm$.001 0.236$\pm$.074 0.299$\pm$.044 0.188$\pm$.028
(Lower the Better) Cot-GAN 1.337$\pm$.068 0.408$\pm$.086 0.980$\pm$.071 1.094$\pm$.079
TimeLDM 0.013$\pm$.005 0.028$\pm$.009 0.028$\pm$.009 0.181$\pm$.029
Diffusion-TS 0.015$\pm$.004 0.004$\pm$.001 0.049$\pm$.008 0.193$\pm$.027
Correlational TimeGAN 0.045$\pm$.010 0.063$\pm$.005 0.210$\pm$.006 0.886$\pm$.039
Score TimeVAE 0.131$\pm$.010 0.095$\pm$.008 0.111$\pm$.020 0.388$\pm$.041
Diffwave 0.022$\pm$.005 0.030$\pm$.020 0.175$\pm$.006 0.579$\pm$.018
DiffTime 0.017$\pm$.004 0.006$\pm$.002 0.067$\pm$.005 0.218$\pm$.031
(Lower the Better) Cot-GAN 0.049$\pm$.010 0.087$\pm$.004 0.249$\pm$.009 1.042$\pm$.007
TimeLDM 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Diffusion-TS 0.006$\pm$.007 0.067$\pm$.015 0.061$\pm$.009 0.008$\pm$.002
Discriminative TimeGAN 0.011$\pm$.008 0.102$\pm$.021 0.114$\pm$.055 0.238$\pm$.068
Score TimeVAE 0.041$\pm$.044 0.145$\pm$.120 0.209$\pm$.058 0.230$\pm$.102
Diffwave 0.017$\pm$.008 0.232$\pm$.061 0.190$\pm$.008 0.203$\pm$.096
DiffTime 0.013$\pm$.006 0.097$\pm$.016 0.100$\pm$.007 0.154$\pm$.045
(Lower the Better) Cot-GAN 0.254$\pm$.137 0.230$\pm$.016 0.325$\pm$.099 0.426$\pm$.022
TimeLDM 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
Diffusion-TS 0.093$\pm$.000 0.036$\pm$.000 0.119$\pm$.002 0.007$\pm$.000
TimeGAN 0.093$\pm$.019 0.038$\pm$.001 0.124$\pm$.001 0.025$\pm$.003
Predictive TimeVAE 0.093$\pm$.000 0.039$\pm$.000 0.126$\pm$.004 0.012$\pm$.002
Score Diffwave 0.093$\pm$.000 0.047$\pm$.000 0.130$\pm$.001 0.013$\pm$.000
DiffTime 0.093$\pm$.000 0.038$\pm$.001 0.121$\pm$.004 0.010$\pm$.001
(Lower the Better) Cot-GAN 0.100$\pm$.000 0.047$\pm$.001 0.129$\pm$.000 0.068$\pm$.009
Original 0.094$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.007$\pm$.001
@@ 4.1 Experimental Setups
Table 2: Main results on multiple time series datasets.The best result in each case is bolded.
@@
Setups. In this paper, all the neural networks used for the encoder, latent diffusion, and decoder of TimeLDM are implemented with PyTorch [34] package. For the well-training across all datasets, we tune the limited hyperparameter, as shown in Table 3. Especially, the batch sizes of VAE and LDM are the same. We proceed with the training in two steps. The first step is to train the $\beta$-VAE to obtain latent space information. The second step involves training a diffusion model in the latent space. In the first step, we optimize our network using Adam with default decay rates. The initial learning rate is $10^{-3}$. In the second step, We optimize our network using Adam with $\left(b_{1},b_{2}\right)=(0.9,0.96)$. The initial learning rate start is $10^{-4}$. The main results are training on an NVIDIA RTX 4080 GPU.
Parameter Sines Stocks ETTh MuJoCo
Attention heads 2 2 2 1
Attention head dimension 16 16 16 32
Encoder layers 1 2 2 2
Decoder layers 2 3 3 3
Batch size 1024 512 1024 512
Hidden dimension of LDM 1024 1024 1024 1024
Table 3: Hyperparameters of VAE and latent diffusion model.
@@
Evaluation Methods. For quantitative analysis, we adopt four different evaluation metrics to evaluate the synthesized time series: (1) Context-Fréchet Inception Distance (Context-FID) score [16] assesses the quality of the synthetic time series samples by calculating the difference between representations of time series that fit into the local context; (2) Correlational score [30] assesses temporal dependencies by calculating the absolute error between the cross-correlation matrices of real and synthetic data. (3) Discriminative score [13] evaluates similarity by employing a classification model to differentiate between original and synthetic data in a supervised setting. (4) Predictive score [13] assesses the utility of synthesized data by training a sequence model post-hoc to predict future temporal vectors using the train-synthesis-and-test-real (TSTR) method. For qualitative analysis, we apply two different data representation methods to evaluate the synthesized time series: (1) t-SNE [31] evaluates synthesized time series by projecting both original and synthetic data into a 2-dimensional space; (2) Kernel density estimation is to draw data distributions to check the alignment between original and synthetic data.
4.2 Unconditional Time Series Generation
@@
Main Results. We follow the previous setup in TimeGAN [13] to analyze the performance of models on the benchmark datasets mentioned above. The quantitative evaluation results of 24-length time series generation, which represents the most common basis of comparison in existing works, are listed in Table 2. As can be seen, TimeLDM shows better performance in most cases than Diffusion-TS which is the previous state-of-the-art model. In particular, compared with the Context-FID score and Discriminative score, our TimeLDM achieves state-of-the-art performance on all benchmarks, with the average improvement of 3.4$\times$ in the Context-FID score and 3.8$\times$ in the Discriminative score, respectively. Figure LABEL:fig:ldm shows the qualitative evaluation results of t-SNE, where red dots represent real data instances and blue dots represent generated data samples in all plots. The figure reveals that our methods have better performance with better overlap between the generated data and with original data. Figure 1 presents the distribution alignment of original data and synthetic information. The TimeLDM aligns better with the original data than Diffusion-TS.
Further Analysis. To further confirm the scalability of our TimeLDM, we evaluate the impact of the different time-series lengths on the generative models for unconditional time series. We examine ETTh data with two different lengths, 64 and 128. For these experiments, we keep all the same hyperparameters with the same metrics to assess the generation quality of different methods. The quantitative results are reported in Table 4. As we can see, our proposed TimeLDM can achieve better performance in most evaluation metrics. Especially on the Context-FID score and Discriminative score, TimeLDM realizes significant state-of-the-art performance. The qualitative results are depicted in Figure 3 and 4. Following the previous methods [4], we randomly generate 3000 sequences to project original and synthetic data into a 2-dimensional space. Our TimeLDM shows better alignment than Diffusion-TS with the original data.
Figure 3: Distribution of our TimeLDM and Diffusion-TS on ETTh datasets with different time series lengths. The blue solid line is for original data, and the yellow dotted line is for synthetic data. TimeLDM aligns better with the original ETTh data.
Figure 4: t-SNE visualization on ETTh data. Red is for original data, and blue represents synthetic data. More overlap between the blue and red dots indicates better matching of the data distributions from the generated to the original samples. Our TimeLDM demonstrates a more accurate overlap with the original data across two different lengths of the ETTh dataset.
@@
Table 4: Further evaluaction on long-term ETTh dataset generation. The best result in each case is bolded.
4.3 Ablation Study
@@
In this part, we first assess the adaptive $\beta$ with the fixed values in the VAE model. Then, we analyze the effectiveness of the reconstruction loss function of VAE. We compare the loss function with its three variants: (1) w/o FFT term during training, (2) w/o $\mathcal{L}_{1}$ norm term during training, (3) w/o $\mathcal{L}_{2}$ norm term during training. The ablation study across all the benchmarks presents the results in Table 5 and Table 6, respectively.
@@
The effect of adaptive $\beta$. We evaluate the adaptive weighting coefficient $\beta$ in the VAE model. Table 5 presents the results of adaptive $\beta$ and constant $\beta$ on the aforementioned datasets. As can be seen, the difference in performance between the Sines and Stocks benchmarks is not significant. While there is a significant performance disparity between the ETTh and MuJoCo benchmarks. The adaptive $\beta$ improve the effectiveness of TimeLDM, remarkably. This emphasizes the superior performance demonstrated by the adaptive $\beta$ approach in training the VAE model.
Metric $\beta$ Sines Stocks ETTh MuJoCo
Discriminative Adaptive 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Score $10^{-2}$ 0.004$\pm$.004 0.015$\pm$.015 0.034$\pm$.019 0.095$\pm$.007
(Lower the Better) $10^{-5}$ 0.004$\pm$.004 0.020$\pm$.019 0.093$\pm$.004 0.010$\pm$.007
Adaptive 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
Predictive $10^{-2}$ 0.093$\pm$.000 0.037$\pm$.000 0.122$\pm$.005 0.010$\pm$.001
Score $10^{-5}$ 0.093$\pm$.000 0.038$\pm$.000 0.123$\pm$.005 0.009$\pm$.002
(Lower the Better) Original 0.094$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.007$\pm$.001
@@
Table 5: Ablation study for the adaptive $\beta$, which balances the reconstruction loss and KL loss. The best result in each case is bolded.
The effect of reconstruction loss. We evaluate the effectiveness of reconstruction loss term in the VAE model. Table 6 presents the results of three variants on the aforementioned datasets. As can be seen, the performance gap among them for the Predictive score is negligible, while for the Discriminative score, as the Figure LABEL:fig:ABS, the FFT-loss term shows a crucial role in improving the capability of TimeLDM.
Metric Methods Sines Stocks ETTh MuJoCo
TimeLDM 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Discriminative w/o FFT 0.008$\pm$.003 0.022$\pm$.028 0.013$\pm$.009 0.010$\pm$.009
Score w/o $\mathcal{L}_{1}$ 0.007$\pm$.005 0.020$\pm$.015 0.014$\pm$.014 0.004$\pm$.006
(Lower the Better) w/o $\mathcal{L}_{2}$ 0.006$\pm$.005 0.014$\pm$.010 0.010$\pm$.008 0.005$\pm$.005
TimeLDM 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
w/o FFT 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.006 0.008$\pm$.002
Predictive w/o $\mathcal{L}_{1}$ 0.093$\pm$.000 0.037$\pm$.000 0.121$\pm$.004 0.008$\pm$.001
Score w/o $\mathcal{L}_{2}$ 0.093$\pm$.000 0.037$\pm$.000 0.121$\pm$.006 0.008$\pm$.002
(Lower the Better) Original 0.094$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.007$\pm$.001
@@
Table 6: Ablation study for VAE reconstruction loss function. The best result in each case is bolded.
Metric Methods Sines Stocks ETTh MuJoCo
Discriminative Score w/o VAE 0.019 $\pm$.015 0.062$\pm$0.029 0.011$\pm$.007 0.023$\pm$.028
(Lower the Better) TimeLDM 0.006$\pm$.005 0.017$\pm$.011 0.009$\pm$.003 0.005$\pm$.004
Predictive Score w/o VAE 0.095 $\pm$.000 0.039 $\pm$.000 0.124 $\pm$.003 0.011$\pm$.001
(Lower the Better) TimeLDM 0.093$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.008$\pm$.001
Table 7: Ablation study for VAE representation. The best result in each case is bolded.
The effect of Latent Space Information. As shown in the Table 7, the w/o VAE method means the removal of Variational Autoencoder(VAE), thereby directly generating time series information in the data space. Obviously, the absence of VAE for mapping feature data into latent space results in significant performance drops.
5 Conclusion

### v2

@@ III-D Training and Sampling
IV EXPERIMENTS
We evaluate TimeLDM for time series generation with five different benchmarks, covering simulated and real-world datasets. Our framework demonstrates better performance than existing methods, both qualitatively and quantitatively. Further analysis across various lengths of time series data confirms the robustness of TimeLDM.
Metric Methods Sines MuJoCo
TimeLDM 0.004$\pm$.001 0.006$\pm$.000
Diffusion-TS 0.006$\pm$.000 0.013$\pm$.001
Context-FID TimeGAN 0.101$\pm$.014 0.563$\pm$.052
Score TimeVAE 0.307$\pm$.060 0.251$\pm$.015
DiffTime 0.006$\pm$.001 0.188$\pm$.028
(Lower the Better) Cot-GAN 1.337$\pm$.068 1.094$\pm$.079
TimeLDM 0.013$\pm$.005 0.189$\pm$.029
Diffusion-TS 0.015$\pm$.004 0.193$\pm$.027
Correlational TimeGAN 0.045$\pm$.010 0.886$\pm$.039
Score TimeVAE 0.131$\pm$.010 0.388$\pm$.041
DiffTime 0.017$\pm$.004 0.218$\pm$.031
(Lower the Better) Cot-GAN 0.049$\pm$.010 1.042$\pm$.007
TimeLDM 0.006$\pm$.005 0.004$\pm$.004
Diffusion-TS 0.006$\pm$.007 0.008$\pm$.002
Discriminative TimeGAN 0.011$\pm$.008 0.238$\pm$.068
Score TimeVAE 0.041$\pm$.044 0.230$\pm$.102
DiffTime 0.013$\pm$.006 0.154$\pm$.045
(Lower the Better) Cot-GAN 0.254$\pm$.137 0.426$\pm$.022
TimeLDM 0.093$\pm$.000 0.007$\pm$.000
Diffusion-TS 0.093$\pm$.000 0.007$\pm$.000
TimeGAN 0.093$\pm$.019 0.025$\pm$.003
Predictive TimeVAE 0.093$\pm$.000 0.012$\pm$.002
Score DiffTime 0.093$\pm$.000 0.010$\pm$.001
(Lower the Better) Cot-GAN 0.100$\pm$.000 0.068$\pm$.009
Original 0.094$\pm$.001 0.007$\pm$.001
@@ IV EXPERIMENTS
TABLE II: Main results on simulated time series datasets. The best result in each case is bolded.
Figure 3: Visualizations of the simulated MuJoCo dataset, synthesized by TimeLDM, Diffusion-TS and TimeVAE.
IV-A Experimental Setups
Datasets. We utilize five different datasets to evaluate our model, including: Sine is a simulated dataset with 5 features in sinusoidal sequence, and each feature has independent frequencies and phases [9]; MuJoCo is the multivariate physics simulation time series data with 14 features [35]; Stocks is the Google stock price information from 2004 to 2019, presented daily information and includes 6 features [7]; ETTh is built from electricity transformers on 15 minutes basis, including load and oil temperature from July 2016 to July 2018 [36]; fMRI serves as a benchmark for causal discovery, featuring simulations that realistically mimic blood-oxygen-level-dependent time series [37].
Baseline. We compare our TimeLDM against five unconditional time series generation methods, including Diffusion-based architectures (Diffusion-TS [14] and DiffTime [13]), GAN-based models (TimeGAN [9] and Cot-GAN [12]), and VAE-based approach (TimeVAE [7]).
Metric Methods Stocks ETTh fMRI
TimeLDM 0.032$\pm$.007 0.034$\pm$.003 0.139$\pm$.025
Diffusion-TS 0.147$\pm$.025 0.116$\pm$.010 0.105$\pm$.006
Context-FID TimeGAN 0.103$\pm$.013 0.300$\pm$.013 1.292$\pm$.218
Score TimeVAE 0.215$\pm$.035 0.805$\pm$.186 14.449$\pm$.969
DiffTime 0.236$\pm$.074 0.299$\pm$.044 0.340$\pm$.015
(Lower the Better) Cot-GAN 0.408$\pm$.086 0.980$\pm$.071 7.813$\pm$.550
TimeLDM 0.028$\pm$.009 0.028$\pm$.009 1.036$\pm$.025
Diffusion-TS 0.004$\pm$.001 0.049$\pm$.008 1.411$\pm$.042
Correlational TimeGAN 0.063$\pm$.005 0.210$\pm$.006 23.502$\pm$.039
Score TimeVAE 0.095$\pm$.008 0.111$\pm$.020 17.296$\pm$.526
DiffTime 0.006$\pm$.002 0.067$\pm$.005 1.501$\pm$.048
(Lower the Better) Cot-GAN 0.087$\pm$.004 0.249$\pm$.009 26.824$\pm$.449
TimeLDM 0.017$\pm$.011 0.009$\pm$.003 0.102$\pm$.020
Diffusion-TS 0.067$\pm$.015 0.061$\pm$.009 0.167$\pm$.023
Discriminative TimeGAN 0.102$\pm$.021 0.114$\pm$.055 0.484$\pm$.042
Score TimeVAE 0.145$\pm$.120 0.209$\pm$.058 0.476$\pm$.044
DiffTime 0.097$\pm$.016 0.100$\pm$.007 0.245$\pm$.051
(Lower the Better) Cot-GAN 0.230$\pm$.016 0.325$\pm$.099 0.492$\pm$.018
TimeLDM 0.037$\pm$.000 0.118$\pm$.007 0.099$\pm$.000
Diffusion-TS 0.036$\pm$.000 0.119$\pm$.002 0.099$\pm$.000
TimeGAN 0.038$\pm$.001 0.124$\pm$.001 0.126$\pm$.002
Predictive TimeVAE 0.039$\pm$.000 0.126$\pm$.004 0.113$\pm$.003
Score DiffTime 0.038$\pm$.001 0.121$\pm$.004 0.100$\pm$.000
(Lower the Better) Cot-GAN 0.047$\pm$.001 0.129$\pm$.000 0.185$\pm$.003
Original 0.036$\pm$.001 0.121$\pm$.005 0.090$\pm$.001
TABLE III: Main results on real-world time series datasets. The best result in each case is bolded.
@@ IV-A Experimental Setups
Setups. In this paper, all the neural networks of TimeLDM are implemented with PyTorch [38] package. For the well-training across all datasets, we tune the limited hyperparameter, as shown in Table I. We proceed with the training in two steps. The first step is to train the $\beta$-VAE to obtain latent space information. The second step involves training a diffusion model in the latent space. In the first step, we optimize our network using Adam with default decay rates. The initial learning rate is $10^{-3}$. In the second step, We optimize our network using the Adam optimizer with the first and second moment decay rates set to $0.9$ and $0.96$ respectively. The initial learning rate start is $10^{-4}$. The main results are training on an NVIDIA RTX 4080 GPU.
Figure 4: Visualizations of the real-world ETTh dataset, synthesized by TimeLDM, Diffusion-TS, and TimeVAE.
@@
Evaluation Methods. For quantitative analysis, we adopt four different evaluation metrics to evaluate the synthesized time series: (1) Context-Fréchet Inception Distance (Context-FID) score assesses the quality of the synthetic time series samples by calculating the difference between representations of time series that fit into the local context [11]; (2) Correlational score assesses temporal dependencies by calculating the absolute error between the cross-correlation matrices of real and synthetic data [39]; (3) Discriminative score evaluates similarity by employing a classification model to differentiate between original and synthetic data in a supervised setting [9]; (4) Predictive score assesses the utility of synthesized data by training a sequence model post-hoc to predict future temporal vectors using the train-synthesis-and-test-real (TSTR) method [9]. For qualitative analysis, we apply two different data representation methods to evaluate the synthesized time series: (1) t-SNE evaluates synthesized time series by projecting both original and synthetic data into a two-dimensional space [40]; (2) Kernel density estimation is to draw data distributions to check the alignment between original and synthetic data.
Figure 5: Examples of generating time series from the fMRI dataset. Our approach yields the closest results to the original training data.
IV-B Unconditional Time Series Generation
@@
Main Results. We follow the previous setup in TimeGAN [9] to analyze the performance of models on the benchmark datasets mentioned above. The quantitative evaluation results of 24-length time series generation, which represents the most common comparison in existing works, are listed in Table II and III. As can be seen, TimeLDM achieves state-of-the-art results on the simulated benchmarks. Compared with the discriminative score, TimeLDM achieves an average improvement of 55% over Diffusion-TS [14] in all benchmarks. Figure 3 and 4 show the qualitative evaluation results of t-SNE and Kernel density. For the t-SNE analysis, where a greater overlap of blue and red dots shows a better distributional similarity between the generated data and original data. Figure 3 and 4 of t-SNE reveals that our methods have better overlap between the generated data and original data. The Kernel density presents the distribution alignment of original data and synthetic information. Based on the figure, our TimeLDM aligns better with the original data than Diffusion-TS and TimeVAE. We also present the generating time series from the fMRI dataset in Figure 5. Compared to the Diffusion-TS [14] and TimeVAE [7], TimeLDM generates time series that more closely resemble the original training set, while TimeVAE [7] struggles to learn features from the fMRI dataset.
@@
TABLE IV: Further evaluation on long-term ETTh dataset generation. The best result in each case is bolded.
Figure 6: Visualizations of the real-world ETTh dataset, synthesized by TimeLDM and Diffusion-TS with 64 and 128-time series lengths.
Further Analysis. To further confirm the scalability of our TimeLDM, we evaluate the impact of the different time-series lengths on the generative models for unconditional time series. We examine ETTh data with two different lengths, 64 and 128. For these experiments, we keep all the same hyperparameters with the same metrics to assess the generation quality of different methods. The quantitative results are reported in Table IV. As we can see, our proposed TimeLDM can achieve better performance in most evaluation metrics. Especially on the Context-FID score and Discriminative score, TimeLDM realizes significant state-of-the-art performance with 80% and 50% improvement over Diffusion-TS [14]. The qualitative results are depicted in Figure 5. Our TimeLDM shows better alignment than Diffusion-TS with the original data.
IV-C Ablation Study
@@
In this part, we first assess the adaptive $\beta$ with the fixed values ($\beta_{\max}$, $\beta_{\min}$) in the VAE model. Then, we analyze the effectiveness of the reconstruction loss function of VAE. We compare the loss function with its three variants: (1) w/o FFT loss term during training, (2) w/o $\mathcal{L}_{1}$ norm term during training, (3) w/o $\mathcal{L}_{2}$ norm term during training. The ablation study across all the benchmarks presents the results in Table V and Table VI, respectively.
@@
The effect of adaptive $\beta$. We evaluate the adaptive weighting coefficient $\beta$ in the VAE model. Table V presents the results of adaptive $\beta$ and constant values ($\beta_{\max}$, $\beta_{\min}$) on the aforementioned datasets. As can be seen, the difference in performance between the Sines and Stocks benchmarks is insignificant. At the same time, there is a significant performance disparity between the ETTh and MuJoCo benchmarks. The adaptive $\beta$ improve the effectiveness of TimeLDM, remarkably. This emphasizes the superior performance demonstrated by the adaptive $\beta$ approach in training the VAE model.
Metric $\beta$ Sines MuJoCo Stocks ETTh fMRI
Discriminative Adaptive 0.006$\pm$.005 0.004$\pm$.004 0.017$\pm$.011 0.009$\pm$.003 0.102$\pm$.020
Score $10^{-2}$ 0.004$\pm$.004 0.258$\pm$.023 0.015$\pm$.015 0.034$\pm$.019 0.496$\pm$.003
(Lower the Better) $10^{-5}$ 0.004$\pm$.004 0.007$\pm$.008 0.020$\pm$.019 0.093$\pm$.004 0.216$\pm$.021
Adaptive 0.093$\pm$.000 0.007$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.099$\pm$.000
Predictive $10^{-2}$ 0.093$\pm$.000 0.013$\pm$.001 0.037$\pm$.000 0.122$\pm$.005 0.100$\pm$.000
Score $10^{-5}$ 0.093$\pm$.000 0.007$\pm$.001 0.038$\pm$.000 0.123$\pm$.005 0.100$\pm$.000
(Lower the Better) Original 0.094$\pm$.001 0.007$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.090$\pm$.001
@@
TABLE V: Ablation study for the adaptive $\beta$, which balances the reconstruction loss and KL loss. The best result in each case is bolded.
Metric Methods Sines MuJoCo Stocks ETTh fMRI
TimeLDM 0.006$\pm$.005 0.004$\pm$.004 0.017$\pm$.011 0.009$\pm$.003 0.102$\pm$.020
Discriminative w/o FFT 0.008$\pm$.003 0.005$\pm$.002 0.022$\pm$.028 0.013$\pm$.009 0.115$\pm$.019
Score w/o $\mathcal{L}_{1}$ 0.007$\pm$.005 0.008$\pm$.004 0.020$\pm$.015 0.014$\pm$.014 0.107$\pm$.019
(Lower the Better) w/o $\mathcal{L}_{2}$ 0.006$\pm$.005 0.007$\pm$.006 0.014$\pm$.010 0.010$\pm$.008 0.129$\pm$.014
TimeLDM 0.093$\pm$.000 0.007$\pm$.000 0.037$\pm$.000 0.118$\pm$.007 0.099$\pm$.000
w/o FFT 0.093$\pm$.000 0.008$\pm$.002 0.037$\pm$.000 0.118$\pm$.006 0.099$\pm$.000
Predictive w/o $\mathcal{L}_{1}$ 0.093$\pm$.000 0.008$\pm$.002 0.037$\pm$.000 0.121$\pm$.004 0.099$\pm$.000
Score w/o $\mathcal{L}_{2}$ 0.093$\pm$.000 0.007$\pm$.001 0.037$\pm$.000 0.121$\pm$.006 0.099$\pm$.000
(Lower the Better) Original 0.094$\pm$.001 0.007$\pm$.001 0.036$\pm$.001 0.121$\pm$.005 0.090$\pm$.001
@@
TABLE VI: Ablation study for VAE reconstruction loss function. The best result in each case is bolded.
The effect of reconstruction loss. We evaluate the effectiveness of reconstruction loss term in the VAE model. Table VI presents the results of three variants on the aforementioned datasets. As can be seen, the performance gap among them for the Predictive score is negligible, the FFT loss term and $\mathcal{L}_{1}$ norm term show a crucial role in improving the capability of TimeLDM for the Discriminative score.
V CONCLUSIONS